Tommy’s age is the sum of the ages of Carmen and Dennis.

Carmen’s age is 2 more than the sum of Dennis and Mark.

Dennis’ age is 4 times Mark’s age. The sum of all four ages is

42. How old is Mark?


1
Expert's answer
2021-09-22T22:54:11-0400

let Mark's "=M"

Therefore, Dennis' age "=4M"

Carmen's age "= [2+(4M+M)]=2+5M"

Tommy's age"=(2+5M)+(4M)=2+9M"

Now,

"M+4M+2+5M+2+9M=42"

"19M=38"

"M=\frac{38}{19}=2"

Therefore, Mark is 2 years old.